# Augmenting Multimodal LLMs with Self-Reflective Tokens for Knowledge-based Visual Question Answering

### v1

@@ Abstract
Multimodal LLMs (MLLMs) are the natural extension of large language models to handle multimodal inputs, combining text and image data. They have recently garnered attention due to their capability to address complex tasks involving both modalities. However, their effectiveness is limited to the knowledge acquired during training, which restricts their practical utility. In this work, we introduce a novel method to enhance the adaptability of MLLMs by integrating external knowledge sources. Our proposed model, Reflective LLaVA (ReflectiVA), utilizes reflective tokens to dynamically determine the need for external knowledge and predict the relevance of information retrieved from an external database. Tokens are trained following a two-stage two-model training recipe. This ultimately enables the MLLM to manage external knowledge while preserving fluency and performance on tasks where external knowledge is not needed. Through our experiments, we demonstrate the efficacy of ReflectiVA for knowledge-based visual question answering, highlighting its superior performance compared to existing methods. Source code and trained models are publicly available at https://github.com/aimagelab/ReflectiVA. Equal contribution.
@@ 2 Related Work
Multimodal Large Language Models. Large Language Models (LLMs) [64, 28, 16] have shown remarkable performance and adaptability across diverse tasks [12, 63], with recent extensions incorporating multimodal capabilities, particularly integrating vision and language [48, 62, 37, 7]. A primary challenge in this domain is effectively combining LLMs with visual features and creating multimodal datasets for robust training. Architectural solutions for vision-to-language integration vary, including single linear projections or MLPs, as in the LLaVA models [43, 44], as well as Q-Former [38, 14] or Perceiver [33] modules to extract fixed-dimensional visual features. Models like Flamingo [1, 3] utilize cross-attention layers to directly integrate multimodal information. For training, multiple stages often leverage image captions for visual-text alignment [18, 10], and specialized datasets are developed for visual instruction tuning [43, 14, 32], with data quality and annotation specificity shown to significantly influence performance [15].
@@ 4.1 Experimental Setup
Datasets. Our experiments are conducted on Encyclopedic-VQA [49] and InfoSeek [11], which contain question-answer pairs linked to documents from an external knowledge base (e.g. Wikipedia). Encyclopedic-VQA consists of 221k pairs associated with 16.7k fine-grained entities (i.e. Wikipedia pages). Questions are divided into single-hop and two-hop types: the former indicates that a single Wikipedia page is required to answer them, while the latter requires a sequential retrieval process across multiple documents. Dataset samples are split into training, validation, and test sets with 1M, 13.6k, and 5.8k items respectively. Experiments are reported on the test set, where single-hop questions correspond to 4.8k samples. The InfoSeek dataset, instead, contains 1.3M image-question pairs associated with around 11k Wikipedia pages. The dataset comprises 934k training, 73k validation, and 348k test samples. Following existing literature [8, 69], experimental results are reported on the validation set which includes questions not contained in the training split and questions associated with unseen entities.
@@
Architectural and Training Details. Our model is based on the LLaVA-v1.5 MLLM [44] with LLaMA-3.1-8B [16] as language model [^4] https://huggingface.co/aimagelab/LLaVA_MORE-llama_3_1-8B-finetuning. It employs CLIP ViT-L/14@336 as visual encoder and an MLP as the vision-to-language connector. For both training phases, we fine-tune the LLaVA architecture to learn how to generate the introduced reflective tokens. To do that, we modify the original vocabulary of the LLaMA-3.1 LLM, replacing the final four reserved special tokens with our custom tokens. During training, we employ a learning rate of $2\times 10^{-5}$ and a global batch size of $128$, updating the weights of both the MLP and LLM.
@@
Table 2: VQA accuracy scores on the Encyclopedic-VQA test set and the InfoSeek validation set, where all results from retrieval-augmented models are reported without considering any re-ranking stage to reorder retrieved documents. Gray color indicates results that are not directly comparable due to different knowledge bases, and the marker $\diamondsuit$ represents our reproductions with different LLMs.
@@ 4.2 Comparison with the State of the Art
Results on Encyclopedic-VQA and InfoSeek. We evaluate our model on the aforementioned datasets, comparing it to various zero-shot LLMs, MLLMs, and retrieval-augmented competitors. Specifically, we report results of three LLMs – Vicuna [12], LLaMA-3, and LLaMA-3.1 [16] – each prompted with both the query question and a description of the query image generated by an image captioning model. Additionally, we assess the performance of BLIP-2 [38], InstructBLIP [14], and LLaVA-v1.5 [44], without external retrieval augmentation and using only the query image and question as input. As direct competitors, we include DPR [35], RORA-VLM [53], Wiki-LLaVA [8], and EchoSight [69], which all leverage external knowledge retrieval. To ensure a fair comparison with the considered methods, for both Wiki-LLaVA and EchoSight we develop a variant based on LLaMA-3.1, employing the same knowledge bases used in our solution.
@@
Table 3: VQA accuracy scores on Encyclopedic-VQA when models are equipped with a document re-ranking component. Gray color indicates results that are not directly comparable due to different knowledge bases, and the marker $\diamondsuit$ represents our reproductions.
@@
ViQuAE S3VQA
Model LLM F1 EM GPT-4
LLaVA-v1.5 [44] Vicuna-7B 15.1 26.6 23.9
LLaVA-v1.5 [44] LLaMA-3.1-8B 15.0 25.6 24.4
Wiki-LLaVA (E-VQA) [8] LLaMA-3.1-8B 10.5 16.7 22.7
Wiki-LLaVA (InfoSeek) [8] LLaMA-3.1-8B 12.7 21.8 21.8
ReflectiVA (w/o KB) LLaMA-3.1-8B 16.6 27.6 26.9
23.2 38.1 29.3
ReflectiVA (Ours) LLaMA-3.1-8B (52.0%) (16.8%)
@@ 4.3 Ablation Studies and Analyses
Effectiveness of Two-Stage Training. We analyze the impact of our two-stage, two-model training strategy. Specifically, we assess the effectiveness of the in-article reflective model, which is trained using both positive and negative passages drawn from the same Wikipedia pages, on Encyclopedic-VQA and InfoSeek. Table 5 (top) demonstrates that training only the first model alone does not yield high accuracy scores. Notably, the complete model (i.e. ReflectiVA) consistently achieves higher accuracy.
Effectiveness of Reflective Tokens. To evaluate the effectiveness of the proposed reflective tokens, we conduct ablation studies that isolate the contributions of both retrieval and relevance tokens. We design three variants of the inference pipeline: (i) without retrieval from the knowledge base (i.e. always enforcing <NORET> during generation), (ii) without assessing the relevance of retrieved passages, where two random passages are selected from each of the top-5 documents, and (iii) always enforcing the <RET> token while using the standard <REL>/<NOREL> pipeline. Results are summarized in the top part of Table 5 for Encyclopedic-VQA and InfoSeek. As shown, consistently performing retrieval causes only a minor performance degradation, given that both datasets are designed to require external knowledge. In contrast, omitting the proposed relevance tokens leads to a substantial accuracy drop (e.g. from 35.5 to 23.6 on Encyclopedic-VQA and from 40.1 to 31.4 on InfoSeek), underscoring the critical role of identifying relevant passages before answer generation. Furthermore, the lowest scores occur when retrieval from the external knowledge base is entirely bypassed, emphasizing the need for retrieval.
@@
Results on Other Knowledge-based Datasets. We also validate the generalization capabilities of ReflectiVA to zero-shot settings which always require knowledge retrieval. Specifically, we report the results on two additional knowledge-based VQA datasets, i.e. ViQuAE [34] and S3VQA [27]. For the former, we include performance in terms of F1 score and exact matching, while for the latter we ask GPT-4 to compute a score between 0 and 100 of the correctness of the given answers, following [9, 67, 51] [^7] Details on datasets and evaluation are reported in the supplementary.. From Table 6, it can be seen that even in these challenging settings, ReflectiVA achieves the best performance, demonstrating the usefulness of predicting reflective tokens.
@@ Acknowledgments
We acknowledge the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources and support. This work has been conducted under a research grant co-funded by Altilia s.r.l., and supported by the PNRRM4C2 project “FAIR - Future Artificial Intelligence Research”, funded by the European Commission, and by the PNRR project “Italian Strengthening of Esfri RI Resilience” (ITSERR) funded by the European Union - NextGenerationEU (CUP B53C22001770006).
@@ B.1 Datasets
Additional Knowledge-based VQA Datasets. In the main paper (cf. Table 6), we perform zero-shot experiments on two additional knowledge-based VQA datasets, namely ViQuAE [34] and S3VQA [27]. Specifically, we report experiments on the ViQuAE test set, composed of 3,317 image-questions pairs, and on the S3VQA validation set, that contains 750 samples. To facilitate the experiments, for both datasets, we employ the 2M knowledge base of Encyclopedic-VQA, performing image-to-image retrieval with EVA-CLIP as the retrieval model. To compute the results, for ViQuAE samples we follow the official evaluation protocol which includes F1 and exact match scores. Instead, given the high level of difficulty of the questions contained in S3VQA and the absence of an official evaluation protocol, we evaluate generated answers using GPT-4. In detail, following recent literature [9], we prompt GPT-4 with the question and the image description extracted with BLIP-2 [38] [^9] https://huggingface.co/Salesforce/blip2-flan-t5-xl, and ask the model to evaluate the alignment between the predicted and ground-truth answers with a score from 0 to 100, where 0 indicates no alignment and 100 indicates a perfect alignment between the two answers101010 Specifically, the prompt we used in our evaluation is: You are trying to evaluate the alignment between a predicted answer and a ground-truth answer for a given question-image pair. To do this, consider the context provided by the question itself and the caption of the query image. # Question: {question} # Image Caption: {caption} # Ground-truth Answer: {ground-truth answer} # Predicted Answer: {predicted answer} You have to determine the alignment between the predicted answer and the ground-truth on a scale from 0 to 100, where 0 indicates no alignment and 100 indicates perfect alignment. Your response should be in JSON format, outputting a list where each element is a dictionary representing a candidate with: ‘‘score’’: a numeric value between 0 and 100 indicating the alignment level, ‘‘reason’’: a string explaining the rationale for the given score. .
@@ B.2 Additional Implementation Details
For simplicity, here we include a single textual passage. However, as detailed in Sec. 3.1 of the main paper, the final answer is generated using the set of passages that have been judged relevant. As shown, we also introduce two additional special tokens, <paragraph> and </paragraph>, to mark the beginning and end of each passage. These tokens are incorporated into the original LLaMA-3.1 vocabulary, as described for the proposed reflective tokens (cf. Sec. 4.1).
@@ C.1 Further Ablation Studies and Analyses
Accuracy scores are presented in Table 11, where we also evaluate the impact of applying our strategy after the first or second stage of LLaVA training. LLaVA-v1.5 follows a two-stage training process: the first stage pre-trains on image-caption pairs to align image features with the LLM textual space, while the second stage focuses on enhancing multimodal conversational capabilities. We therefore analyze the optimal point in this pipeline to incorporate our strategy for learning new special tokens. Firstly, we can observe that accuracy scores are consistently higher than 85% for all reflective tokens, highlighting the ability of the model to accurately predict when retrieval is necessary and whether the retrieved passages are relevant to the query. Furthermore, applying our training strategy after the second stage of LLaVA training generally yields the best results.
@@
Table 11: Accuracy scores achieved by ReflectiVA in predicting the proposed reflective tokens.
@@
C.2 Does ReflectiVA Integrate Built-In Re-Ranking Capabilities?
Q: Terry park became the pirates spring what ballpark?
Refer to caption Ground-truth: Training home ReflectiVA (Ours):
Training ground
Q: In what cuisine is this vegetable used?
Refer to caption Ground-truth: Southeast Asia ReflectiVA (Ours):
Indian cuisine
Q: What is the fee to see this gate from the outside?
Refer to caption Ground-truth: No admission fee ReflectiVA (Ours):
Free
Q: What is the objective of this object?
Engine tuning
Q: What is the magnitude of the velocity of this
animal kilometer per hour?
Refer to caption Ground-truth: 120 ReflectiVA (Ours):
150
Q: What is the conservation status of this bird?
Refer to caption Ground-truth: Endangered ReflectiVA (Ours):
Vulnerable
Appendix D Qualitative Results

### v2

@@ Abstract
Multimodal LLMs (MLLMs) are the natural extension of large language models to handle multimodal inputs, combining text and image data. They have recently garnered attention due to their capability to address complex tasks involving both modalities. However, their effectiveness is limited to the knowledge acquired during training, which restricts their practical utility. In this work, we introduce a novel method to enhance the adaptability of MLLMs by integrating external knowledge sources. Our proposed model, Reflective LLaVA (ReflectiVA), utilizes reflective tokens to dynamically determine the need for external knowledge and predict the relevance of information retrieved from an external database. Tokens are trained following a two-stage two-model training recipe. This ultimately enables the MLLM to manage external knowledge while preserving fluency and performance on tasks where external knowledge is not needed. Through our experiments, we demonstrate the efficacy of ReflectiVA for knowledge-based visual question answering, highlighting its superior performance compared to existing methods. Source code and trained models are publicly available at https://aimagelab.github.io/ReflectiVA. Equal contribution.
@@ 2 Related Work
Multimodal Large Language Models. Large Language Models (LLMs) [66, 30, 18] have shown remarkable performance and adaptability across diverse tasks [13, 65], with recent extensions incorporating multimodal capabilities, particularly integrating vision and language [50, 64, 39, 8, 15]. A primary challenge in this domain is effectively combining LLMs with visual features and creating multimodal datasets for robust training. Architectural solutions for vision-to-language integration vary, including single linear projections or MLPs, as in the LLaVA models [45, 46], as well as Q-Former [40, 16] or Perceiver [35] modules to extract fixed-dimensional visual features. Models like Flamingo [2, 4] utilize cross-attention layers to directly integrate multimodal information. For training, multiple stages often leverage image captions for visual-text alignment [20, 11], and specialized datasets are developed for visual instruction tuning [45, 16, 34], with data quality and annotation specificity shown to significantly influence performance [17].
@@ 4.1 Experimental Setup
Datasets. Our experiments are conducted on Encyclopedic-VQA [51] and InfoSeek [12], which contain question-answer pairs linked to documents from an external knowledge base (e.g. Wikipedia). Encyclopedic-VQA consists of 221k pairs, each paired with up to 5 images and associated with 16.7k fine-grained entities (i.e. Wikipedia pages). Questions are divided into single-hop and two-hop types: the former indicates that a single Wikipedia page is required to answer them, while the latter requires a sequential retrieval process across multiple documents. Dataset samples are split into training, validation, and test sets with 1M, 13.6k, and 5.8k items respectively. Experiments are reported on the test set, where single-hop questions correspond to 4.8k samples. The InfoSeek dataset, instead, contains 1.3M image-question pairs associated with around 11k Wikipedia pages. The dataset comprises 934k training, 73k validation, and 348k test samples. Following existing literature [9, 71], experimental results are reported on the validation set which includes questions not contained in the training split and questions associated with unseen entities.
@@
Architectural and Training Details. Our model is based on the LLaVA-v1.5 MLLM [46, 15] with LLaMA-3.1-8B [18] as language model [^4] https://huggingface.co/aimagelab/LLaVA_MORE-llama_3_1-8B-finetuning. It employs CLIP ViT-L/14@336 as visual encoder and an MLP as the vision-to-language connector. For both training phases, we fine-tune the LLaVA architecture to learn how to generate the introduced reflective tokens. To do that, we modify the original vocabulary of the LLaMA-3.1 LLM, replacing the final four reserved special tokens with our custom tokens. During training, we employ a learning rate of $2\times 10^{-5}$ and a global batch size of $128$, updating the weights of both the MLP and LLM. The training is stopped based on the loss value on a separate validation split extracted from the training dataset.
@@
Table 2: VQA accuracy scores on the Encyclopedic-VQA test set and the InfoSeek validation set, where all results from retrieval-augmented models are reported without considering any re-ranking stage to reorder retrieved documents. $\dagger$ indicates results that are not directly comparable due to different knowledge bases, and the marker $\diamondsuit$ represents our reproductions with different LLMs.
@@ 4.2 Comparison with the State of the Art
Results on Encyclopedic-VQA and InfoSeek. We evaluate our model on the aforementioned datasets, comparing it to various zero-shot LLMs, MLLMs, and retrieval-augmented competitors. Specifically, we report results of four LLMs – Vicuna [13], LLaMA-3, LLaMA-3.1 [18] and GPT-4 [1] – each prompted with both the query question and a description of the query image generated by an image captioning model. Additionally, we assess the performance of BLIP-2 [40], InstructBLIP [16], LLaVA-v1.5 [46] and GPT-4V [1] without external retrieval augmentation and using only the query image and question as input. As direct competitors, we include DPR [37], RORA-VLM [55], Wiki-LLaVA [9], and EchoSight [71], which all leverage external knowledge retrieval. To ensure a fair comparison with the considered methods, for both Wiki-LLaVA and EchoSight we develop a variant based on LLaMA-3.1, employing the same knowledge bases used in our solution.
@@
Table 3: VQA accuracy scores on Encyclopedic-VQA when models are equipped with a document re-ranking component. $\dagger$ indicates results that are not directly comparable due to different knowledge bases, and the marker $\diamondsuit$ represents our reproductions.
@@
Single model (w/ special tokens from LLaMA-3.1) 5 30.7 27.3 28.2 28.5
@@ 4.3 Ablation Studies and Analyses
Effectiveness of Two-Stage Training. We analyze the impact of our two-stage, two-model training strategy. Specifically, we assess the effectiveness of the in-article reflective model, which is trained using both positive and negative passages drawn from the same Wikipedia pages, on Encyclopedic-VQA and InfoSeek. Table 5 (top) demonstrates that training only the first model alone does not yield high accuracy scores. Likewise, training a single model on full-dataset annotations generated by open-source LLMs such as LLaMA-3.1 fails to achieve competitive performance. In constrast, the complete model (i.e. ReflectiVA) consistently achieves higher accuracy.
Effectiveness of Reflective Tokens. To evaluate the effectiveness of the proposed reflective tokens, we conduct ablation studies that isolate the contributions of both retrieval and relevance tokens. We design four variants of the inference pipeline: (i) always enforcing the <RET> token while using the standard <REL>/<NOREL> pipeline, (ii) without using relevance tokens, instead selecting the top passages retrieved by the Contriever [27] model, (iii) without assessing the relevance of retrieved passages, where two random passages are selected from each of the top-5 documents, and (iv) without retrieval from the knowledge base (i.e. always enforcing <NORET> during generation). Results are summarized in the top part of Table 5 for Encyclopedic-VQA and InfoSeek. As shown, consistently performing retrieval causes only a minor performance degradation, given that both datasets are designed to require external knowledge. In contrast, omitting the proposed relevance tokens leads to a substantial accuracy drop (e.g. from 35.5 to 23.6 on Encyclopedic-VQA and from 40.1 to 31.4 on InfoSeek), underscoring the critical role of identifying relevant passages before answer generation. Furthermore, the lowest scores occur when retrieval from the external knowledge base is entirely bypassed, emphasizing the need for retrieval. Also, replacing the <REL> and <NOREL> tokens with the top-2 passages retrieved by Contriever results in significantly lower performance on both datasets, underscoring the benefit of explicit relevance modeling over direct retrieval.
@@ Acknowledgments
We acknowledge the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources and support. This work has been conducted under a research grant co-funded by Altilia s.r.l., and supported by the PNRRM4C2 project “FAIR - Future Artificial Intelligence Research”, funded by the European Commission, by the PNRR project “Italian Strengthening of Esfri RI Resilience” (ITSERR), funded by the European Union - NextGenerationEU (CUP B53C22001770006), and by the “Fortissimo Plus” (FFplus) project, funded by the European High-Performance Computing Joint Undertaking (No. 101163317). We would also like to thank Davide Caffagni and Sara Sarto for their valuable support.
@@ B.1 Datasets
Additional Knowledge-based VQA Datasets. In addition to the datasets employed in the main paper, we perform zero-shot experiments on two additional knowledge-based VQA datasets, namely ViQuAE [36] and S3VQA [29]. Specifically, we report experiments on the ViQuAE test set, composed of 3,317 image-questions pairs, and on the S3VQA validation set, that contains 750 samples. To facilitate the experiments, for both datasets, we employ the 2M knowledge base of Encyclopedic-VQA, performing image-to-image retrieval with EVA-CLIP as the retrieval model. To compute the results, for ViQuAE samples we follow the official evaluation protocol which includes F1 and exact match scores. Instead, given the high level of difficulty of the questions contained in S3VQA and the absence of an official evaluation protocol, we evaluate generated answers using GPT-4. In detail, following recent literature [10, 69, 53], we prompt GPT-4 with the question and the image description extracted with BLIP-2 [40] [^8] https://huggingface.co/Salesforce/blip2-flan-t5-xl, and ask the model to evaluate the alignment between the predicted and ground-truth answers with a score from 0 to 100, where 0 indicates no alignment and 100 indicates a perfect alignment between the two answers999 Specifically, the prompt we used in our evaluation is: You are trying to evaluate the alignment between a predicted answer and a ground-truth answer for a given question-image pair. To do this, consider the context provided by the question itself and the caption of the query image. # Question: {question} # Image Caption: {caption} # Ground-truth Answer: {ground-truth answer} # Predicted Answer: {predicted answer} You have to determine the alignment between the predicted answer and the ground-truth on a scale from 0 to 100, where 0 indicates no alignment and 100 indicates perfect alignment. Your response should be in JSON format, outputting a list where each element is a dictionary representing a candidate with: ‘‘score’’: a numeric value between 0 and 100 indicating the alignment level, ‘‘reason’’: a string explaining the rationale for the given score. .
@@ B.2 Additional Implementation Details
For simplicity, here we include a single textual passage. However, as detailed in Sec. 3.1 of the main paper, the final answer is generated using the set of passages that have been judged relevant. As shown, we also introduce two additional special tokens, <paragraph> and </paragraph>, to mark the beginning and end of each passage. These tokens are incorporated into the original LLaMA-3.1 vocabulary, as described for the proposed reflective tokens (Sec. 4.1).
@@ B.3 Baselines and Competitor Reproductions
ViQuAE S3VQA
Model LLM F1 EM GPT-4
LLaVA-v1.5 [46] Vicuna-7B 15.1 26.6 23.9
LLaVA-v1.5 [46] LLaMA-3.1-8B 15.0 25.6 24.4
Wiki-LLaVA (E-VQA) [9] LLaMA-3.1-8B 10.5 16.7 22.7
Wiki-LLaVA (InfoSeek) [9] LLaMA-3.1-8B 12.7 21.8 21.8
ReflectiVA (w/o KB) LLaMA-3.1-8B 16.6 27.6 26.9
23.2 38.1 29.3
ReflectiVA (Ours) LLaMA-3.1-8B (52.0%) (16.8%)
@@ C.1 Further Ablation Studies and Analyses
Results on Other Knowledge-based Datasets. In addition to the results on the Encyclopedic-VQA and InfoSeek datasets, we also validate the generalization capabilities of ReflectiVA to zero-shot settings which always require knowledge retrieval. Specifically, we report the results on two additional knowledge-based VQA datasets, i.e. ViQuAE [36] and S3VQA [29]. From Table 9, it can be seen that even in these challenging settings, ReflectiVA achieves the best results, outperforming competitors by a significant margin and demonstrating the usefulness of predicting reflective tokens.
@@
Table 11: Accuracy scores achieved by ReflectiVA in predicting the proposed reflective tokens, compared with the performance of GPT-4 and GPT-4V.
Accuracy scores are shown in Table 11, where we compare the performance of zero-shot models, such as GPT-4 and GPT-4V, with the results obtained by applying our strategy after either the first or second stage of LLaVA training. LLaVA-v1.5 follows a two-stage training process: the first stage pre-trains on image-caption pairs to align image features with the LLM textual space, while the second stage focuses on enhancing multimodal conversational capabilities. We therefore analyze the optimal point in this pipeline to incorporate our strategy for learning new special tokens. As it can be seen, accuracy scores are consistently higher than 85% for all reflective tokens, highlighting the ability of the model to accurately predict when retrieval is necessary and whether the retrieved passages are relevant to the query. Also, applying our training strategy after the second stage of LLaVA training generally yields the best results.
C.2 Does ReflectiVA Integrate Built-In Re-Ranking Capabilities?
Appendix D Qualitative Results
@@ Appendix E Limitations and Failure Cases
Q: Terry park became the pirates spring what ballpark?
Refer to caption Ground-truth: Training home ReflectiVA (Ours):
Training ground
Q: In what cuisine is this vegetable used?
Refer to caption Ground-truth: Southeast Asia ReflectiVA (Ours):
Indian cuisine
Q: What is the fee to see this gate from the outside?
Refer to caption Ground-truth: No admission fee ReflectiVA (Ours):
Free
Q: What is the objective of this object?
Engine tuning
Q: What is the magnitude of the velocity of this
animal kilometer per hour?
Refer to caption Ground-truth: 120 ReflectiVA (Ours):
150
Q: What is the conservation status of this bird?
Refer to caption Ground-truth: Endangered ReflectiVA (Ours):
Vulnerable